# Thermodynamic favourability of 1T phase over 1H phase in group III metal monochalcogenide zigzag nanoribbons

### v2

@@ CONCLUSION
Employing DFT-based calculations, we thoroughly examined the width-dependent structural and electronic properties of hydrogen passivated and bare MX ZNRs. The remarkable finding of this study is the demonstration of the existing crossover between the 1T and 1H phases. Contrary to 2D monolayers, where 1H is the most thermodynamically favorable phase, in MX ZNRs, the 1T phase is more favorable than 1H, below experimentally achievable critical widths ranging from 23 nm in GaS up to 34 nm in InSe ZNRs. The 1T-UP-ZNRs possess a nearly flat top valance band due to localized edge states, and band gaps that do not show substantial width-dependent variation. In 1T-P-ZNRs, the top valence band retains a Mexican-hat shape akin to their 2D monolayer counterparts, and band gaps exhibit a downward trend toward 2D monolayer values, creating a tunable platform for thermoelectric and optoelectronic applications. Another noteworthy observation is that, unlike 1H ZNRs, there is no built-in potential between the edges in 1T ZNRs. Lastly, positive hydrogenation energies along with nearly flat top valence bands make InSe 1T-UP-ZNRs potential candidates for spintronic and thermoelectric applications. Our results lay the foundation for future experimental investigations.
Acknowledgments
@@
E.A., A.M. and S.J. acknowledge support from the TÜBİTAK-ANAS project number 120N296. H.S. acknowledges support from the Air Force Office of Scientific Research (AFOSR, Award No. FA9550-21-1-0261). The calculations were performed at TÜBİTAK ULAKBİM, High Performance and Grid Computing Center (TRUBA resources) and National Center for High-Performance Computing of Turkey (UHeM) under grant number 1007742020.
@@
Figure S2: Calculated formation energies of MX NRs plotted against ribbon width. Solid Lines are guides for the eye.
@@
Figure S6: Spin-polarized band structures of (a) passivated 1H, (b) passivated 1T, (c) unpassivated 1H and (d) unpassivated 1T $N=8$ GaS ZNRs.
Figure S7: Spin-polarized band structures of (a) passivated 1H, (b) passivated 1T, (c) unpassivated 1H and (d) unpassivated 1T $N=8$ GaSe ZNRs.
Figure S8: Partial charge distribution of the HOMO and LUMO bands of (a) unpassivated 1T and (d) passivated 1T $N=8$ InSe ZNRs.

### v3

@@ 1.3 Conclusion
Employing DFT-based calculations, we thoroughly examined the width-dependent structural and electronic properties of hydrogen passivated and bare MX ZNRs. The remarkable finding of this study is the demonstration of the existing crossover between the 1T and 1H phases. Contrary to 2D monolayers, where 1H is the most thermodynamically favorable phase, in MX ZNRs, due to competition between built-in electric field and reconstruction processes, the non-polar 1T phase becomes more favorable than the polar 1H phase below the critical widths spanning from 23 nm in GaS up to 34 nm in InSe ZNRs, a range achievable by state-of-the-art fabrication methods. The 1T-UP-ZNRs possess a nearly flat top valence band due to localized edge states, and band gaps that do not show substantial width-dependent variation. In 1T-P-ZNRs, the top valence band retains a Mexican-hat shape akin to their 2D monolayer counterparts, and band gaps exhibit a downward trend toward 2D monolayer values, creating a tunable platform for thermoelectric and optoelectronic applications. Another noteworthy observation is that, positive hydrogenation energies along with nearly flat top valence bands make InSe 1T-UP-ZNRs potential candidates for spintronic and thermoelectric applications. Our results lay the foundation for future experimental and theoretical investigations.
{acknowledgement}
@@
E.A., A.M. and S.J. acknowledge support from the TÜBİTAK project number 124F108. H.S. acknowledges support from the Air Force Office of Scientific Research (AFOSR, Award No. FA9550-21-1-0261). The calculations were performed at TÜBİTAK ULAKBİM, High Performance and Grid Computing Center (TRUBA resources) and National Center for High-Performance Computing of Turkey (UHeM) under grant number 1007742020.
{suppinfo}
Lattice constants of MX 2D and ZNR structures, formation energies of MX ZNRs, bandgaps of 1T MX ZNRs, planar average of the local electrostatic potential of GaS and GaSe ZNRs for N=8{N}=8, spin-polarized band structures of GaS and GaSe ZNRs for N=8{N}=8; InSe 1T-UP-ZNRs and 1T-P-ZNRs for N=4,5,6,7,9,10{N}=4,5,6,7,9,10; InSe 1H and 1T 2D $1\times 8$ supercell and ZNRs for N=8{N}=8.
@@
Figure S2: Calculated formation energies of MX NRs plotted against ribbon width. Solid Lines are guides to the eye.
@@
Figure S6: Spin-polarized band structures of (a) unpassivated 1H, (b) unpassivated 1T, (c) passivated 1H and (d) passivated 1T $N=8$ GaS ZNRs.
Figure S7: Spin-polarized band structures of (a) unpassivated 1H, (b) unpassivated 1T, (c) passivated 1H and (d) passivated 1T $N=8$ GaSe ZNRs.
Figure S8: Spin-polarized band structures of InSe 1T-UP-ZNRs for N=4,5,6,7,9,10{N}=4,5,6,7,9,10.
Figure S9: Spin-polarized band structures of InSe 1T-P-ZNRs for N=4,5,6,7,9,10{N}=4,5,6,7,9,10.
Figure S10: Spin-polarized band structures of InSe a) 2D 1H $1\times 8$ supercell, b) 1H-P-ZNR8, c) 1H-UP-ZNR8, d) 2D 1T $1\times 8$ supercell, e) 1T-P-ZNR8, and f) 1T-UP-ZNR8.